# Impact of cavities on the detection of quadratically coupled ultra-light dark matter

### v1

@@ III.2 Spherical Cavity
Where $|k|>k_{c}$, the cavity can have a significant effect on the scalar field profile. In the case $\alpha>0$ we observe that both the constant and linear components of the field appear to be rapidly suppressed as $\alpha$ grows. This is corroborated by algebraic expressions for these coefficients in Eqs. (28), (29), the denominators of which grow exponentially with $k/k_{c}$. For the case $\alpha<0$, the behaviours of $A_{0}$ and $A_{1}$ are less simply characterised. Generally, they still follow a decaying trend, though this is power law decay rather than exponential suppression, and thus considerably slower than the $\alpha>0$ case. This is a consequence of the hyperbolic terms in Eqs. (28), (29) becoming oscillatory for negative $\alpha$. Additionally, for negative $\alpha$ there are values of $k$ for which the denominators of the coefficients $A_{0},\leavevmode\nobreak\ A_{1}$ are zero, leading to values of $\alpha$ for which the field inside the cavity is enhanced above the background, as can be seen in Fig. 2. From Eqs. (28), (29), we see $A_{0}$ is explicitly singular where
$\tan(|k|(R_{2}-R_{1}))=\frac{1}{|k|R_{1}}\leavevmode\nobreak\ ,$ (30)
@@
$\tan(|k|(R_{2}-R_{1}))=\frac{3|k|R_{1}}{(|k|R_{1})^{2}-3}\leavevmode\nobreak\ .$ (31)
Whilst we expect enhancement to be physical around these points, caution should be taken since the necessary condition $|\alpha\phi^{2}|\ll 1$ may not hold for all $m,\leavevmode\nobreak\ \alpha$ in their neighbourhoods. It should be noted that Fig. 2 does not exactly show the number, nor the singular nature of these divergences due to limited resolution. Importantly, as cavity wall density, and wall thickness are increased, the threshold value $|\alpha|=k_{c}^{2}/\rho_{c}$ decreases, and hence cavity effects on the field become more significant at weaker couplings.
@@ III.3 Cylindrical Cavity
?figurename? 3: Absolute value of $\varphi$ at the centre (blue) of an aluminium cylindrical cavity with inner radius $R=0.1$ m and wall thickness $\delta R=\delta L=1$ cm. The average field value inside the cavity is shown in orange. The top row corresponds to $\alpha>0$, while the bottom row to $\alpha<0$. The black dashed (dotted) lines mark the analytical result for an infinite cylinder (sphere).
@@
$\nabla^{2}\varphi(r,z)=\alpha\rho(r,z)\varphi(r,z)\enspace,$ (32)
@@
?figurename? 4: Field profile for $\alpha$ corresponding to the first three peaks, $\alpha=(1.41,1.75,2.27)\times 10^{-12}\leavevmode\nobreak\ \mathrm{GeV}^{-2}$ for the long (top), and $\alpha=(2.83,15.0,24.5)\times 10^{-12}\leavevmode\nobreak\ \mathrm{GeV}^{-2}$ for the short (bottom) cavities. Notice that the range of field values differs in each panel. The black line shows the position of the cavity walls.
@@ III.4 Impact on Existing Constraints
Precise modelling of each experiment is beyond the scope of this work. Here we consider a hypothetical, but illustrative, scenario in which each experiment is contained within the spherically symmetric cavity discussed in Sec. III.2, with inner radius $R_{1}=10\mbox{\leavevmode\nobreak\ cm}$, outer radius $R_{2}=11\mbox{\leavevmode\nobreak\ cm}$, and density $\rho=2700\mbox{\leavevmode\nobreak\ kg m}^{-3}$. The revised constraints are shown, orange shaded region, in Fig. 5, as well as the original constraints, dotted line, and regions of validity of the quadratic model, outside the grey hashed region. These regions of validity are determined by requiring: $|d_{e}\kappa^{2}\phi^{2}|,\leavevmode\nobreak\ |(d_{\hat{m}}-d_{g})\kappa^{2}\phi^{2}|,\leavevmode\nobreak\ |(d_{m_{e}}-d_{g})\kappa^{2}\phi^{2}|\ll 1$, for panels (a), (b) and (c) respectively, analogous to the condition $|\alpha\phi^{2}|\ll 1$.
@@
From Fig. 5, we clearly observe the impact of the cavity in the $\alpha>0$ (or: $d_{e},\leavevmode\nobreak\ (d_{\hat{m}}-d_{g}),\leavevmode\nobreak\ (d_{m_{e}}-d_{g})>0$) case for strong couplings. In this regime, we observe significant departure from original constraints, rendering large regions of parameter space potentially unconstrainable in a laboratory or satellite experiment. Additionally, whilst the boundary of the original constraints lies well outside the region in which the quadratic model of interactions is invalid, cavity effects cause our re-mapped values to intersect with this region. Constraints here cannot be developed without a more comprehensive model of scalar interactions, which is beyond the scope of this work.
@@
For $\alpha<0$ (or: $d_{e},\leavevmode\nobreak\ (d_{\hat{m}}-d_{g}),\leavevmode\nobreak\ (d_{m_{e}}-d_{g})<0$), the authors of Ref. [26] exclude a large region of parameter space above the first divergence in $\varphi_{\oplus}$, on the basis of the breakdown of the quadratic model. Given that the solution for the field profile around the Earth (Eq. (18)) diverges only for specific values of $\alpha$, marking the full region as unconstrainable may be a more conservative approach than is needed, although simulating the behaviour of the field in this region is expected to be challenging. Shown in Fig. 5 is our own attempt at plotting a region of validity for the quadratic model, allowing for divergences in the Earth’s field. Again, we can show neither the true number, nor height of these divergences due to plotting resolution. The apparent density of these, particularly at negative values of the dilaton coefficients, is a consequence of of the logarithmic scale of the plot, given that the divergences are linearly spaced. Modifications to any constraints in the $\alpha<0$ panels of Fig. 5 case, arising from cavity effects, are expected to appear in the region below the blue dashed lines - marking approximately where $|\alpha|=k_{c}^{2}/\rho_{c}$.
@@
?figurename? 5: An example of possible modifications to existing constraints on ultra-light quadratically coupled dark matter. We reproduce the constraints of Ref. [26] (black dotted line) derived from tests of the universality of free-fall, and atomic clock measurements. Constrains are plotted in terms of the scalar mass $m$, and the coupling strength in terms of the “maximum reach” dillaton parameters on the left vertical axis. On the right vertical axis we show the equivilent values for our universal coupling $\alpha$, calculated using dilaton charges $Q_{i}$ for iron (see Sec. II.1, and for greater detail Refs. [22, 23]). These reference values may experience an order of magnitude variation depending on the material chosen. Orange regions show the updated parameter space constraints if each experiment was contained inside a spherical cavity with the same properties as described in Fig. 2. This is only illustrative of the effects of a cavity on the constraints, and full reevaluation of the constraints requires careful modelling of each experimental environment. Dark hatched regions depict where the quadratic model breaks down. In light green are regions unconstrained by Ref. [26] for the same reason. Finally, the blue dashed lines approximately correspond to $|k|=k_{c}$ for the cavity solutions, and hence where the cavity enters the regime of strong coupling to the field.
@@ IV Differential 5th Forces Acting on Cavities
We have so far considered cavities as an experimental environment, focussing on the enhancement and suppression of the field within the cavity, and the potential impact this might have on measurements therein. We now consider the evaluation of $5^{\rm th}$ forces acting on spherical cavities themselves, and in particular if changes in the geometry of the cavity, keeping the mass and external radius fixed, leads to detectable differences in the force on the cavities. This is motivated by the fact that test masses of similar geometry have been used in laboratory based $5^{\rm th}$ force searches, in particular the Eöt-Wash experiment, which uses approximately spherical test bodies with identical mass and external dimension, but differing composition [66, 67]. The MICROSCOPE [68, 69] satellite experiment similarly looks for anomalous differential acceleration between cylindrical, equal-mass Titanium and Platinum bodies towards the Earth. Data from both have been used to set constraints on the properties of scalar field dark matter [26]. There are also proposals for an experiment to measure the gravitational force on a Casimir cavity [70]. However, the situation we consider is significantly simplified compared to the experimental environment that would be required for a measurement.
@@ IV.1 Force on a Single Cavity
$\vec{F}_{5}=-\frac{1}{2}\alpha\rho_{c}\int_{V}d^{3}x\leavevmode\nobreak\ \nabla\phi^{2}$ (33)
@@
$\vec{F}_{5}=-\frac{4\pi\rho_{c}\alpha}{3}\left[R_{2}^{2}\left(a_{0}+\frac{B^{out}_{0}}{R_{2}}\right)\left(a_{1}R_{2}+\frac{B^{out}_{1}}{R_{2}^{2}}\right)-R_{1}^{3}A^{in}_{0}A^{in}_{1}\right]\phi_{H}^{2}\leavevmode\nobreak\ \hat{z}\;,$ (34)
@@
$\vec{F}_{5}^{point}=-\alpha M_{c}a_{0}a_{1}\phi_{H}^{2}\leavevmode\nobreak\ .$ (35)
This point mass approximation gives the false impression that two test bodies with identical masses and values for $\alpha$ will experience identical forces. The non-trivial dependence of Eq. (34) on $R_{1},\leavevmode\nobreak\ R_{2}$, shows that this is not necessarily the case. In particular, the point mass approximation fails to accurately describe the force on a body in the strongly coupled limit. The force in Eq. (34) tends asymptotically to a constant limit as $|\alpha|$ is increased:
@@
In the weak coupling regime between and the field, the point mass solution is more successful. For small $|\alpha|$, Eq. (34) takes the form:
@@
$M_{c}=\frac{4\pi}{3}\rho_{c}\left(R_{2}^{3}-R_{1}^{3}\right)\leavevmode\nobreak\ .$ (38)
@@ IV.2 Differential Force Evaluation
We now use Eq. (34) to evaluate the differential Earth-cavity $5^{\rm th}$ force on two spherical cavities labelled $1$ and $2$:
@@
[…]
Figure 6 shows possible constraints imposed by the measurement of the magnitude of the differential $5^{th}$ force to a precision $\sim 100\mbox{ fN}$ using two cavities with the properties shown in Table 1, assuming a universal coupling of the scalar to matter. The figure also shows (in the shaded region), where $|\alpha\phi^{2}|>1$. In this region of the parameter space we can no longer trust that higher order interactions, in the action of Eq. (1), between the dark matter scalar field and matter are small. As before, this region is complex, in the $\alpha<0$, case as a consequence of divergences induced by both the Earth and cavities. For reference, the black dotted line shows the constraints of Ref. [26] on the dilaton coeffiecient $d_{e}$, assuming maximum reach, expressed in terms of the coupling strength $\alpha$, using equations (7) and (9), as in Figure 5.
@@
?tablename? 1: Cavity properties used in the generation of Fig. 6
[…]
Both panels of Fig. 6, for positive and negative $\alpha$, show constraints with similar general behaviours. Both show constrained regions which initially extend to greater $m$ as $|\alpha|$ increases. This derives from a differential force which initally grows with $|\alpha|$. At $|\alpha|=1/(\rho_{\oplus}R_{\oplus}^{2})\sim 10^{-29}$ GeV-2, the trajectory of the boundary of these regions changes subtly - a consequence of $k_{\oplus}$ now satisfying $|k_{\oplus}|R_{\oplus}\gg 1$, and resulting in changes to the background field profile due to the Earth. More significantly, around $|\alpha|\sim|k^{2}_{c}|/{\rho_{c}}$, which is $\mathcal{O}\left(10^{-11}\mbox{ GeV}^{-2}\right)$ for both cavities, the boundary of the constrained regions begins to retract to smaller $m$. This is a consequence of the force on each cavity reaching the $\alpha$ independent limit, outlined in Eq. (36). This limit derives from the finite size of the cavities, and is identical for those in Table 1, due to their identical external radii $R_{2}$. This effect is not captured in constraints which treat test bodies as point masses.
In the $\alpha<0$ case, Fig. 6 additionally demonstrates spikes in the boundary of the constrained region, associated with the diverging field at discrete values of $\alpha$. When $|\alpha|\gtrsim 1/(\rho_{\oplus}R_{\oplus}^{2})$, these are due to divergences in the background profile due to the Earth. When $\alpha\gtrsim k_{c}^{2}/\rho_{c}$, the divergences are additionally induced by the cavity. Not all of these divergences are visible in the Figure due to finite resolution.
We also observe in, Fig. 6, that of the three values of $h$ considered, assuming a height $h=1000\mbox{ km}$ constrains the largest region of parameter space. Indeed, the $5^{\rm th}$ force in Eq. (34), which depends on $h$ via a factor $a_{0}a_{1}$, is maximised with respect to $h$ where:
@@
We additionally highlight that Fig. 6 does not include any suppression, particularly relevant if $\alpha>0$, of the differential force due the to environment of the experiment, akin to the suppression of forces inside the cavity seen in Sec. III.2. We have assumed that the test cavities are positioned in the vicinity of the Earth, but with no other material around them. Surrounding experimental apparatus will, as discussed in Section III.2, result in a more significant suppression of the field at strong couplings.
The significance of the forecasts in Fig. 6 depends on assumptions about whether the dark matter scalar couples universally, or preferentially to certain types of matter. If $\alpha$ is allowed to be composition dependent, as per II.1, we observe that the constraints deriving from purely geometric effects in Fig. 6 are largely contained within regions of already constrained parameter space (e.g. see the black dashed line in the upper panel). Hence, whilst differential forces from geometric effects could be increased by considering larger and denser cavities, we expect that impractical properties would be needed for them to extend existing constraints on models featuring composition dependent couplings. [^4] Our forecasts are derived assuming universal couplings. If there are compositional dependent couplings then this may lead to additional equivalence principle violation depending on the materials of the two cavities.
However, we note that constraints such as those in Ref. [26], illustrated in Fig. 5, are necessarily predicated on the existence of a composition dependent coupling. They are therefore not applicable to models featuring universal $\alpha$. For such models, Fig. 6 demonstrates that an experiment designed to be sensitive to geometrically induced differential $5^{\rm{th}}$ forces could be a viable means to constrain parameter space. “CubeSats” - with similar form factor and mass to the cavities considered here - may present themselves as a means by which to search for such effects. These present two main advantages. Foremost, they may be placed in orbit at high altitudes which, as shown in Fig. 6, may improve the capacity for the experiment to constrain new parameter space. Second, such devices are not contained within further experimental apparatus, and so do not experience the environmental suppression described in Sec.II, to which experiments enclosed inside satellites, for example - MICROSCOPE [68, 69], may be vulnerable. An experiment could be realised in which a pair of CubeSats, of similar mass and scale but differing internal structure, could be set into orbit, from which differential deviation from their trajectories could be studied to place bounds on $5^{\rm{th}}$ forces.

### v2

@@ III.2 Spherical Cavity
Where $|k|>k_{c}$, the cavity can have a significant effect on the scalar field profile. In the case $\alpha>0$ we observe that both the constant and linear components of the field appear to be rapidly suppressed as $\alpha$ grows. This is corroborated by algebraic expressions for these coefficients in Eqs. (28), (29), the denominators of which grow exponentially with $k/k_{c}$. For the case $\alpha<0$, the behaviours of $A_{0}$ and $A_{1}$ are less simply characterised. Generally, they still follow a decaying trend, though this is power law decay rather than exponential suppression, and thus considerably slower than the $\alpha>0$ case. This is a consequence of the hyperbolic terms in Eqs. (28), (29) becoming oscillatory for negative $\alpha$. Additionally, for negative $\alpha$ there are values of $k$ for which the denominators of the coefficients $A_{0},\penalty 10000\ A_{1}$ are zero, leading to values of $\alpha$ for which the field inside the cavity is enhanced above the background, as can be seen in Fig. 3. From Eqs. (28), (29), we see $A_{0}$ is explicitly singular where
$\tan(|k|(R_{2}-R_{1}))=\frac{1}{|k|R_{1}}\penalty 10000\ ,$ (30)
@@
$\tan(|k|(R_{2}-R_{1}))=\frac{3|k|R_{1}}{(|k|R_{1})^{2}-3}\penalty 10000\ .$ (31)
Whilst we expect enhancement to be physical around these points, caution should be taken since the necessary condition $|\alpha\phi^{2}|\ll 1$ may not hold for all $m,\penalty 10000\ \alpha$ in their neighbourhoods. It should be noted that Fig. 3 does not exactly show the number, nor the singular nature of these divergences due to limited resolution. Importantly, as cavity wall density, and wall thickness are increased, the threshold value $|\alpha|=k_{c}^{2}/\rho_{c}$ decreases, and hence cavity effects on the field become more significant at weaker couplings.
@@ III.3 Cylindrical Cavity
Figure 4: Absolute value of $\varphi$ at the centre (blue) of an aluminium cylindrical cavity with inner radius $R=0.1$ m and wall thickness $\delta R=\delta L=1$ cm. The average field value inside the cavity is shown in orange. The top row corresponds to $\alpha>0$, while the bottom row to $\alpha<0$. The black dashed (dotted) lines mark the analytical result for an infinite cylinder (sphere).
@@
$\nabla^{2}\varphi(r,z)=\alpha\rho(r,z)\varphi(r,z)\kern 5.0pt,$ (32)
@@
Figure 5: Field profile for $\alpha$ corresponding to the first three peaks, $\alpha=(1.41,1.75,2.27)\times 10^{-12}\penalty 10000\ \mathrm{GeV}^{-2}$ for the long (top), and $\alpha=(2.83,15.0,24.5)\times 10^{-12}\penalty 10000\ \mathrm{GeV}^{-2}$ for the short (bottom) cavities. Notice that the range of field values differs in each panel. The black line shows the position of the cavity walls.
@@ III.4 Impact on Existing Constraints
Precise modelling of each experiment is beyond the scope of this work. Here we consider a hypothetical, but illustrative, scenario in which each experiment is contained within the spherically symmetric cavity discussed in Sec. III.2, with inner radius $R_{1}=10\mbox{\penalty 10000\ cm}$, outer radius $R_{2}=11\mbox{\penalty 10000\ cm}$, and density $\rho=2700\mbox{\penalty 10000\ kg m}^{-3}$. The revised constraints are shown, orange shaded region, in Fig. 6, as well as the original constraints, dotted line, and regions of validity of the quadratic model, outside the grey hashed region. These regions of validity are determined by requiring: $|d_{e}\kappa^{2}\phi^{2}|,\penalty 10000\ |(d_{\hat{m}}-d_{g})\kappa^{2}\phi^{2}|,\penalty 10000\ |(d_{m_{e}}-d_{g})\kappa^{2}\phi^{2}|\ll 1$, for panels (a), (b) and (c) respectively, analogous to the condition $|\alpha\phi^{2}|\ll 1$.
@@
From Fig. 6, we clearly observe the impact of the cavity in the $\alpha>0$ (or: $d_{e},\penalty 10000\ (d_{\hat{m}}-d_{g}),\penalty 10000\ (d_{m_{e}}-d_{g})>0$) case for strong couplings. In this regime, we observe significant departure from original constraints, rendering large regions of parameter space potentially unconstrainable in a laboratory or satellite experiment. Additionally, whilst the boundary of the original constraints lies well outside the region in which the quadratic model of interactions is invalid, cavity effects cause our re-mapped values to intersect with this region. Constraints here cannot be developed without a more comprehensive model of scalar interactions, which is beyond the scope of this work.
@@
For $\alpha<0$ (or: $d_{e},\penalty 10000\ (d_{\hat{m}}-d_{g}),\penalty 10000\ (d_{m_{e}}-d_{g})<0$), the authors of Ref. [53] exclude a large region of parameter space above the first divergence in $\varphi_{\oplus}$, on the basis of the breakdown of the quadratic model due to divergences in field amplitude. Given that the solution for the field profile around the Earth (Eq. (18)) diverges only for specific values of $\alpha$, marking the full region as unconstrainable may be a more conservative approach than is needed, although simulating the behaviour of the field in this region is expected to be challenging. Shown in Fig. 6 is our own attempt at plotting a region of validity for the quadratic model, incorporating both conditions on the stability of solutions, and the requirement $|\alpha|\phi^{2}<1$ for stable solutions. Our region is ultimately identical for $m<\sqrt{\rho_{\oplus}|\alpha|}$, but is less exclusive for more massive fields. Modifications to any constraints in the $\alpha<0$ panels of Fig. 6 case, arising from cavity effects, are expected to appear in the region below the blue dashed lines - marking approximately where $|\alpha|=k_{c}^{2}/\rho_{c}$.
@@
Figure 6: An example of possible modifications to existing constraints on ultra-light quadratically coupled dark matter. We reproduce the constraints of Ref. [53] (black dotted line) derived from tests of the universality of free-fall, and atomic clock measurements. Constrains are plotted in terms of the scalar mass $m$, and the coupling strength in terms of the “maximum reach” dilaton parameters on the left vertical axis. On the right vertical axis we show the equivilent values for our universal coupling $\alpha$, calculated using dilaton charges $Q_{i}$ for iron (see Sec. II.1, and for greater detail Refs. [31, 32]). These reference values may experience an order of magnitude variation depending on the material chosen. Orange regions show the updated parameter space constraints if each experiment was contained inside a spherical cavity with the same properties as described in Fig. 3. This is only illustrative of the effects of a cavity on the constraints, and full reevaluation of the constraints requires careful modelling of each experimental environment. Dark hatched regions depict where the quadratic model of interactions breaks down: where $|\alpha\phi^{2}|\gtrsim 1$ and, in the case $\alpha<0$, where field solutions around the Earth are expected to be unstable. In light green are regions unconstrained by Ref. [53] for the same reason. Finally, the blue dashed lines approximately correspond to $|k|=k_{c}$ for the cavity solutions, and hence where the cavity enters the regime of strong coupling to the field.
@@ IV Differential 5th Forces Acting on Cavities
We have so far considered cavities as an experimental environment, focussing on the enhancement and suppression of the field within the cavity, and the potential impact this might have on measurements therein. We now consider the evaluation of $5^{\rm th}$ forces acting on spherical cavities themselves. In particular, we ask if changes in the geometry of the cavity, keeping the mass and external radius fixed, leads to detectable differences in the force on the cavities. This is motivated by the fact that test masses of similar geometry have been used in laboratory based $5^{\rm th}$ force searches; for example the Eöt-Wash experiment, which uses approximately spherical test bodies with identical mass and external dimension, but differing composition [67, 74]. The MICROSCOPE [70, 71] satellite experiment additionally looks for anomalous differential acceleration between cylindrical Titanium and Platinum bodies towards the Earth. Data from both have been used to set constraints on the properties of scalar field dark matter [53]. There are also proposals for an experiment to measure the gravitational force on a Casimir cavity [27]. However, the situation we consider is significantly simplified compared to the experimental environment that would be required for a measurement.
@@ IV.1 Force on a Single Cavity
$\vec{F}_{5}=-\frac{1}{2}\alpha\rho_{c}\int_{V}d^{3}x\penalty 10000\ \nabla\phi^{2}\;.$ (33)
@@
$\vec{F}_{5}=-\frac{4\pi\rho_{c}\alpha}{3}\left[R_{2}^{2}\left(a_{0}+\frac{B^{out}_{0}}{R_{2}}\right)\left(a_{1}R_{2}+\frac{B^{out}_{1}}{R_{2}^{2}}\right)-R_{1}^{3}A^{in}_{0}A^{in}_{1}\right]\phi_{H}^{2}\penalty 10000\ \hat{z}\;$, (34)
@@
$\vec{F}_{5}^{point}=-\alpha M_{c}a_{0}a_{1}\phi_{H}^{2}\penalty 10000\ .$ (35)
This point mass approximation gives the false impression that two test bodies with identical masses and values for $\alpha$ will experience identical forces. The non-trivial dependence of Eq. (34) on $R_{1},\penalty 10000\ R_{2}$, shows that this is not necessarily the case. In particular, the point mass approximation fails to accurately describe the force on a body in the strongly coupled limit. The force in Eq. (34) tends asymptotically to a constant limit as $|\alpha|$ is increased:
@@
In the regime of weak coupling between the cavity and the field, the point mass solution is more successful. For small $|\alpha|$, Eq. (34) takes the form:
@@
$M_{c}=\frac{4\pi}{3}\rho_{c}\left(R_{2}^{3}-R_{1}^{3}\right)\penalty 10000\ .$ (38)
@@ IV.2 Differential Force Evaluation
We now use Eq. (34) to evaluate the differential Earth-cavity $5^{\rm th}$ force on two spherical cavities labelled 11 and 22:
@@
Using the approximate description of the Earth as a uniform spherical body, we determine the coefficients $a_{0}$ and $a_{1}$ as in Eq. (25). We make the an illustrative choice to place the cavity at a heights: $h=100\mbox{ m},10\mbox{ km, and }1000\mbox{ km}$ above the Earth’s surface, where in Eq. (25) $r_{c}=R_{\oplus}+h$. In Sec. IV.3 we discuss further the values of $h$ for which the linearised approximation of field boundary conditions is valid. We eliminate the time dependence in Eq. (34), from the background oscillations in $\phi_{H}^{2}$, by averaging over the oscillations, [^3] Ponderomotive effects of such oscillations are considered in Ref. [78]. allowing us to evaluate an order-of-magnitude estimate of the force experienced by the cavity.
Figure 7 shows possible constraints imposed by the measurement of the magnitude of the differential $5^{th}$ force to a precision $\sim 100\mbox{ fN}$ using two cavities with the properties shown in Table 1, assuming a universal coupling of the scalar to matter. Only forecasts for $\alpha>0$ are presented on account of the regions of instability of our quadratically interacting scalar theory for $\alpha<0$, demonstrated in Fig. 6 and discussed in Sec. III.1. The figure also shows (in the shaded region), where $|\alpha\phi^{2}|>1$. In this region of the parameter space we can no longer trust that higher order interactions, in the action of Eq. (1), between the dark matter scalar field and matter are small. For reference, the black dotted line shows the constraints of Ref. [53] on the dilaton coefficient $d_{e}$, assuming maximum reach, and expressed in terms of the coupling strength $\alpha$ using equations (7) and (9), as in Figure 6.
@@
Table 1: Cavity properties used in the generation of Fig. 7
Figure 7: Forecast constraints on the $\alpha>0,m$ parameter space of a light scalar with quadratic couplings to matter, deriving from differential $5^{\rm th}$ forces experienced by cavities in the vicinity of the Earth. The properties of these cavities are drawn from Table 1. Constrained regions presume that no signal is discovered up to a sensitivity $\sim 100\mbox{ fN}$, akin to current high-precision experiments. Constraints are plotted presuming measurements to be taken at a height $h=100\mbox{ m},10\mbox{ km, and }1000\mbox{ km}$ above the surface of the Earth. Blue lines indicate characteristic values of $\alpha$ for the system: The dashed line shows $|\alpha|\sim k_{c}^{2}/{\rho_{c}}$, and the dot-dashed line shows $|\alpha|\sim 1/(\rho_{\oplus}R_{\oplus}^{2})$. For both cavities 11 and 22, the value of $k_{c}$ differs only by a factor of 22, and hence we only show the line for cavity 1 in this figure. For the $\alpha>0$ case, the black dotted line gives approximate constraints on $\alpha$, deriving from Ref. [53] under the “maximal reach” assumption. As discussed in Sec. IV.1 and Sec. III.4, these constraints do not incorporate environmental suppression effects, nor effects deriving from the finite size of test masses where relevant. The black dotted line should not be used as an indicator of pre-existing constraints for universal $\alpha$. Dark hatched regions indicate where the quadratic model breaks down due to the violation of the condition: $|\alpha\phi^{2}|\ll 1$ at some point in space.
Fig. 7 demonstrates constrained regions of parameter space which initially extend to greater $m$ as $\alpha$ increases. This derives from a differential force which initially grows with $\alpha$. At $\alpha=1/(\rho_{\oplus}R_{\oplus}^{2})\sim 10^{-29}$ GeV-2, the trajectory of the boundary of these regions changes subtly - a consequence of $k_{\oplus}$ now satisfying $|k_{\oplus}|R_{\oplus}\gg 1$, and resulting in changes to the background field profile due to the Earth. More significantly, around $\alpha\sim|k^{2}_{c}|/{\rho_{c}}$, which is $\mathcal{O}\left(10^{-11}\mbox{ GeV}^{-2}\right)$ for both cavities, the boundary of the constrained regions begins to retract to smaller $m$. This is a consequence of the force on each cavity reaching the $\alpha$ independent limit, outlined in Eq. (36). This limit derives from the finite size of the cavities, and is identical for those in Table 1, due to their identical external radii $R_{2}$. This effect is not captured in constraints which treat test bodies as point masses.
We also observe in Fig. 7 that, of the three values of $h$ considered, assuming a height $h=1000\mbox{ km}$ constrains the largest region of parameter space. Indeed, the $5^{\rm th}$ force in Eq. (34), which depends on $h$ via the product $a_{0}a_{1}$, is maximised with respect to $h$ where:
@@
We additionally highlight that Fig. 7 does not include any suppression of the differential force due the to environment of the experiment, akin to the suppression of forces inside the cavity seen in Sec. III.2. We have assumed that the test cavities are positioned in the vicinity of the Earth, but with no other material around them. If there is surrounding experimental apparatus, as discussed in Section III.2, we expect a more significant suppression of the field at strong couplings.
The significance of the forecasts in Fig. 7 depends on assumptions about whether the dark matter scalar couples universally, or preferentially to certain types of matter. If $\alpha$ is allowed to be composition dependent, as per II.1, we observe that the constraints deriving from purely geometric effects in Fig. 7 are largely contained within regions of already constrained parameter space (e.g. see the black dashed line in the upper panel). Hence, whilst differential forces from geometric effects could be increased by considering larger cavities, we expect that impractical properties would be needed for them to extend existing constraints on models featuring composition dependent couplings. [^4] Our forecasts are derived assuming universal couplings. If there are compositional dependent couplings then this may lead to additional equivalence principle violation depending on the materials of the two cavities. The described experiment should be a viable means to constrain parameter space if composition dependence is taken into account.
However, we note that constraints such as those in Ref. [53], illustrated in Fig. 6, are necessarily predicated on the existence of a composition dependent coupling. They are therefore not applicable to models featuring universal $\alpha$. For such models, Fig. 7 demonstrates that an experiment designed to be sensitive to geometrically induced differential $5^{\rm{th}}$ forces could be a viable means to constrain parameter space. “CubeSats” - with similar form factor and mass to the cavities considered here - may present themselves as a means by which to search for such effects. These present two main advantages. Foremost, they may be placed in high altitude orbits which, as shown in Fig. 7, may improve the sensitivity of the experiment to new parameter space. Secondly, because the satellites are themselves the test masses of the experiment, and are not contained within further apparatus, they do not experience the environmental suppression described in Sec.III.2. This is in contrast to experiments such as MICROSCOPE [70, 71], in which the relevant test masses are instead enclosed inside the satellite. An experiment could be realised in which a pair of CubeSats, of similar mass and scale but differing internal structure, could be set into orbit, from which differential deviation from their trajectories could be studied to place bounds on $5^{\rm{th}}$ forces.